OpenStudy (anonymous):

Determine the 2 numbers whose difference is 12 and whose product is a maximum.I came up with (x + 6)^2 - 36 But, what do I do now? It is functions, grade 11. Do I need to explain how I got x + 6)^2 - 36?

OpenStudy (amistre64):

well, how DID you come up with it?

OpenStudy (amistre64):

x+y = 12 ; y = 12-x xy = max x(12-x) = max 12x - x^2 = max at its vertex

OpenStudy (amistre64):

whose difference is 12 ... i gotta get better glasses

OpenStudy (amistre64):

x-y = 12 ; y = x-12 xy = max x(x-12) = max x^2 - 12x = max at infinity

OpenStudy (anonymous):

i repeat, maybe it means two numbers whose sum is 12 if the difference is 12 i can make it as large as possible

OpenStudy (anonymous):

100,112 100000, 100012 etc

OpenStudy (amistre64):

id do sum too :)

OpenStudy (anonymous):

I have no idea!

OpenStudy (anonymous):

i am going to bet it means two numbers that add to 12 not "whose difference is 12"

OpenStudy (anonymous):

the answer is (0,9) it's at the back of the book, but how????

OpenStudy (amistre64):

9*0 = 0 not 12

OpenStudy (anonymous):

ok now we really have a problem because it is not the case that the difference or the sum of 0 and 9 is 12!!

OpenStudy (amistre64):

theres a valid thought in there someplace lol

OpenStudy (anonymous):

either there was a typo in the book or this is the answer to a different problem 0 + 9 = 9, 9 - 0 = 9, there is no 12 to be had here

OpenStudy (anonymous):

I'm sorry, it's a wrong question the answer is + and - 6

OpenStudy (anonymous):

produce is a minimum maybe ??

OpenStudy (anonymous):

Itr makes sence now, but in my formula it is just 6, where do we get the - 6?

OpenStudy (anonymous):

nope, says max

OpenStudy (anonymous):

now way that \[6\times -6=- 36\] is a maximum

OpenStudy (anonymous):

#no way

OpenStudy (anonymous):

but even if it say minimum ( there is similar question in the book but with min), how the heck we will get -6 and +6?

OpenStudy (anonymous):

what about 100 and 112 the difference is 6 and the product is much bigger than -36

OpenStudy (anonymous):

This question drives me nuts lol

OpenStudy (anonymous):

i can see why because it makes no sense at all

OpenStudy (anonymous):

minimum represents the direction of the porabola ( opens up )

OpenStudy (anonymous):

we don't really care about it

OpenStudy (anonymous):

we can solve this problem find two numbers whose sum is 12 and whose product is a maximum

OpenStudy (anonymous):

I don't knwo....

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

That's fine. I'll ask the teacher